OpenStudy (anonymous):

Rename these fractions to lowest terms:

OpenStudy (anonymous):

9/18 x 4/8 = 36/165 =

OpenStudy (karatechopper):

36/165 can you try dividing by three?

OpenStudy (anonymous):

7/20 x 5/8 = 35/160 =

OpenStudy (anonymous):

divide 5 into 35 and into 160 to get your simplified fraction

OpenStudy (anonymous):

Answer: 7/32!!!! Yay! Yes!

OpenStudy (anonymous):

56/90 x 45/50 = 2520/4500 =